# Long-Duration Station-Keeping Strategy for Cislunar Spacecraft Formations

### v1

@@ 3.3 Local Toroidal Coordinates About HFEM Trajectories
Local toroidal coordinates (LTCs) offer a convenient way to visualize, represent, and encode equilibrium in relative orbital geometries in restricted three-body problem orbits [27, 17]. As a result of the procedure used to identify finite-duration QPROs about HFEM trajectories, it is possible to leverage LTCs for the duration of the QPRO despite the fact that the underlying reference is not periodic. Let $e_{r}+e_{i}i$ represent the quasi-unimodular eigenvalues and $\bm{w}_{r,0}+\bm{w}_{i,0}i$ represent their associated eigenvectors determined from a HFEM STM $\Phi(t_{1},t_{0})$. The mapping of these eigenvectors to any point in time along the reference trajectory is given by $\bm{w}_{r}(t)=\Phi(t,t_{0})\bm{w}_{r,0}$ and $\bm{w}_{i}(t)=\Phi(t,t_{0})\bm{w}_{i,0}$. Let the eigenvectors be decomposed into position and velocity components with $\bm{w}_{r}=\begin{bmatrix}\bm{r}_{r}^{\intercal},\bm{v}_{r}^{\intercal}\end{bmatrix}^{\intercal}$ and $\bm{w}_{i}=\begin{bmatrix}\bm{r}_{i}^{\intercal},\bm{v}_{i}^{\intercal}\end{bmatrix}^{\intercal}$. Local toroidal coordinates are defined as $\bm{\zeta}\triangleq\begin{bmatrix}\bm{\zeta}_{r}^{\intercal},\bm{\zeta}_{v}^{\intercal}\end{bmatrix}^{\intercal}$, $\bm{\zeta}_{r}\triangleq[\alpha,\beta,h]^{\intercal}$, $\bm{\zeta}_{v}\triangleq[\dot{\alpha},\dot{\beta},\dot{h}]^{\intercal}$ and can be related to a relative position and velocity with respect to the reference with
$\begin{gathered}\delta\bm{r}(t)=T(t)\bm{\zeta}_{r},\quad T(t)=\begin{bmatrix}\bm{r}_{r}(t)&\bm{r}_{i}(t)&\hat{\bm{n}}(t)\end{bmatrix},\quad\hat{\bm{n}}(t)\parallel\bm{r}_{r}(t)\times\bm{r}_{i}(t)\\ \delta\bm{v}(t)=\dot{T}(t)\bm{\zeta}_{r}+T(t)\bm{\zeta}_{v},\quad\dot{T}(t)=\begin{bmatrix}\bm{v}_{r}(t)&\bm{v}_{i}(t)&\dot{\hat{\bm{n}}}(t)\end{bmatrix}.\end{gathered}$ (10)
If the LTCs of a spacecraft are such that if $h=0$ and $\bm{\zeta}_{v}=\bm{0}$, then the point lies on the surface of a locally invariant torus. Moreover, the coordinates of $\alpha$ and $\beta$ formation geometry, as is visualized in Fig. 6.
Figure 6: Formation configuration on an osculating invariant circle. Spacecraft can be separated by phase angle ($\theta_{1}$) or radius ($K$), and equivalently represented in local toroidal coordinates ($\alpha$,$\beta$).
@@ 4.1 Absolute Control
To compute absolute station-keeping maneuvers that keep the formation in the vicinity of a reference, this work makes use of an MPC scheme similar to the one proposed in Ref. 14. However, the scheme proposed here instead computes just one maneuver (instead of two) that targets a position ellipsoid at perilune crossing six revolutions downstream along the reference. This is akin to the popularly employed x-axis crossing scheme [12, 31], which involves computing one maneuver to target the state of the reference trajectory at six revolutions downstream, which has been found to produce favorable $\Delta V$ expenditures. Given a reference trajectory $\bm{\phi}(t)$, an initial condition $\bm{x}_{0}$ and initial time $t_{0}$, the $\Delta V_{abs}$ is solved for with
@@
[…]
where $B=\begin{bmatrix}\bm{0}&I\end{bmatrix}^{\intercal}$, $t_{f}$ is the time corresponding to perilune crossing at six revolutions downstream from the current time, and $\bm{f}$ is a function representing the HFEM equations of motion of the system. Given the nonlinear nature of the optimization problem, it is solved using sequential convex programming (SCP) by enforcing a trust region about the solution variables, augmenting the dynamics constraint with a slack variable, and penalizing the $l_{2}$ norm of the slack in the objective. The only solution variables in the SCP problem are the initial condition dynamic slack variable $\bm{\nu}\in\mathbb{R}^{6}$, which is augmented to the objective with an $l_{2}$-norm penalty, and the $\Delta V_{abs}\in\mathbb{R}^{3}$, for which a trust region of 1 cm/s is enforced. The final position excursion ellipsoid is set to $\epsilon=$ 25 km. The optimization routine should be theoretically always feasible [14]; and an appropriately applied SCP solution routine robustly and reliably converges [32].
@@
Fig. 8 shows the performance of the proposed absolute station-keeping scheme about the reference trajectory in Fig. 1 simulated with a ground truth including solar radiation pressure and with realistic navigation and control errors. Commanded maneuvers with magnitude below 0.15 cm/s are not executed, in line with existing studies of absolute station-keeping [31, 12].
@@ 4.2 Relative Control
Given the inadequacies of lower fidelity models in relative control tasks over long durations, relative dynamical motion is generated directly about the reference HFEM trajectory to more accurately represent the relative motion. Since state transition matrices are used to compute the HFEM trajectory and the HFEM QPROs, this relative motion is readily available and could even be stored upon or uplinked to a spacecraft to allow for autonomous relative control. Since dispersions from a reference relative trajectory are typically small in a station-keeping scenario, a simple and effective procedure for computing corrective impulsive maneuvers is to simply target a reference position along the QPRO at some later time [17]. Concretely, to compute a relative maneuver $\Delta V_{rel}$ to deliver the $k$th spacecraft to a desired position $\delta\bm{r}_{f,k}$ (e.g., a reference position along a QPRO) at time $t_{f}$ from some initial state $\delta\bm{x}_{0,k}$ at $t_{0}$ simply requires solving
@@
Here, $\Phi_{rv}$ and $\Phi_{rx}$ denotes the upper right block and right block of the STM respectively. Under ideal scenarios, the relative state will converge to the reference equilibrium in just two impulses. It is efficient to compute, scalable to multiple agents, and independent of the choice of dynamics. Though the control law makes no consideration for fuel expenditure, relative motion is typically so sensitive to changes in velocity in the station-keeping scenario that the commanded $\Delta V$ is usually only on the order of millimeters per second.
@@ 4.3 Safety Concept
A particular challenge in cislunar formation control is that there exists no known analytical condition which guarantees passive safety of spacecraft in a formation. In Keplerian orbits, passive safety is guaranteed by maintaining perpendicularity between spacecraft relative positions and the direction of the unstable mode. An analytical condition, known as $e-i$ vector separation, has been derived, which when satisfied provides long duration guarantees of safe spacecraft separation under perturbations [33]. In contrast, there does not exist an equally simple and strong notion of passive safety for cislunar orbits. In fact, from a QPRO equilibrium with kilometer level separation, a change in velocity on the order of millimeters per second can deliver a spacecraft to an unsafe configuration in just the span of weeks. In the case in which a spacecraft misses a maneuver, collisions can potentially occur within days or hours. This phenomenon is attributable to the existence of stable and unstable modes of the time-varying eigensystem, whose directions are constantly changing over the span of an orbit.
@@
Collision risk is usually not considered in studies of cislunar formation flight, though solutions that have been proposed include passively-safe sequential convex programming (SCP-DRIFT), whereby nonlinear optimization routines determine a sequence of maneuvers that guarantees passive safety, and maneuver planning leveraging local-toroidal elements (SCP-QPRIT), which allows for elegant suppression of stable and unstable modes [21]. SCP is computationally expensive and scales poorly with agent count, which makes it impractical for long-duration station-keeping. Additionally, LTCs, which provide weaker guarantees of safety, are not amenable to the inclusion of absolute station-keeping inputs, which are typically spanned by unstable or stable modes.
@@
The safety concept proposed in this work leverages an optimization-based control barrier function (CBF) to provide safety guarantees to the formation. This approach takes advantage of the rarity of collision events, provides recursive safety, and allows for decoupled computation of the absolute and relative control inputs. CBFs typically involve using optimization protocols to “filter” a nominal control input to be safe [34]. Though not often applied to spacecraft and typically considered for systems with continuous control, CBFs and similar procedures have been employed for impulsively and continuously controlled spacecraft in Refs. 35 and 36 for ensuring safety in reinforcement learning routines. The variant of CBF proposed here involves enforcing stochastic keep out zone constraints in a similar manner to passively safe optimal control schemes [37] and represents one of the first applications of safety filtering to cislunar formation flight.
@@
Let $t_{CO,max}$ and $t_{CO,min}$ be the maximum and minimum cutoff time to a spacecraft’s propulsion system following a random missed maneuver event at time $t_{0}$. It is assumed that following some cutoff time in this interval, the propulsion system will be guaranteed to function at the next available control time. Under these conditions, it is possible to “filter” the desired control input of all agents at $t_{0}$ by enforcing keep out ellipses over the evolution of the agents over the time interval. Particularly, the mean state and covariance of an agent can be propagated from some initial estimated mean and covariance provided by the navigation filter $\delta\bm{x}_{0,k}\sim\mathcal{N}(\bm{\mu}_{0,k},\Sigma_{k,0})$ and a desired control input $\Delta V_{k}$ such that
$\begin{gathered}\delta\bm{x}_{k}(t)\sim\mathcal{N}(\bm{\mu}_{k}(t),\Sigma_{k}(t)),\\ \bm{\mu}_{k}(t)=\Phi(t,t_{0})(\bm{\mu}_{k,0}+B\Delta V_{k}),\\ \Sigma_{k}(t)=\Phi(t,t_{0})(\Sigma_{0,k}+BG_{exe,k}G^{\intercal}_{exe,k}B^{\intercal})\Phi^{\intercal}(t,t_{0})+Q(t,t_{0}),\\ \dot{Q}(t,t_{0})=A(t)Q(t,t_{0})+Q(t,t_{0})A^{\intercal}(t)+G_{rel}G_{rel}^{\intercal},\quad Q(t_{0},t_{0})=\bm{0}.\end{gathered}$ (13)
@@
Here, $G_{exe,k}$ is an estimate of the maneuver execution error associated with $\Delta V_{k}$ as determined from Gate’s model [38], $G_{rel}$ is the covariance associated with the perturbative acceleration to the relative states, and $Q(t,t_{0})$ is the integration of these relative perturbative accelerations from $t_{0}$ to $t$. The state can also be propagated in the case of missed maneuvers $\delta\bm{x}_{m,k}\sim\mathcal{N}(\bm{\mu}_{m,k},\Sigma_{m,k})$ simply by setting $\Delta V_{k}=0$. To enforce passive safety, it is necessary to ensure that the distributions of each agent are non-overlapping to within some probability for all time until the next available control input. The approach taken here is to conservatively enforce that the $3\sigma$ ellipsoids are more than a distance $D_{min}$ apart [39], producing a non-convex quadratic constraint. Let $t_{1}$ denote the first control input time after $t_{0}+t_{CO,max}$ and $t_{2}$ denote the first control input time after $t_{0}+2t_{CO,max}$. Then, $\{\Delta V_{safe,k}\}_{k=1}^{M}$ can be solved for from a desired $\{\Delta V_{tot,k}\}_{k=1}^{M}$ with
$\min_{\{\Delta V_{safe,k}\}_{k=1}^{M}}\quad$ $\sum_{k=1}^{M}\|\Delta V_{safe,k}-\Delta V_{tot,k}\|_{2}$ (14a)
s.t. $\left(\mu_{k}(t)-\mu_{j}(t)\right)^{\intercal}\left(3\Sigma_{k}(t)+3\Sigma_{j}(t)+D_{min}^{2}\right)^{-1}\left(\mu_{k}(t)-\mu_{j}(t)\right)\geq 1,$
$\quad\forall t\in[t_{0},t_{2}],\forall k<j$ (14b)
$\left(\mu_{k}(t)-\mu_{m,j}(t)\right)^{\intercal}\left(3\Sigma_{k}(t)+3\Sigma_{m,j}(t)+D_{min}^{2}\right)^{-1}\left(\mu_{k}(t)-\mu_{m,j}(t)\right)\geq 1,$
$\quad\forall t\in[t_{0},t_{1}],\forall k\neq j$ (14c)
By enforcing no collisions over twice the maximum propulsion cutoff time for the nominal spacecraft trajectories, the filtered maneuver profile implicitly guarantees that safety is still guaranteed in the case of more than one spacecraft encountering MMEs at the next control input time. This endows the CBF with recursive safety under the assumption that a spacecraft will be able to execute a maneuver following the cutoff period following an MME.
@@
Input $\{\bm{\mu}_{0,k}\}$, $\{\Sigma_{0,k}\}$, $t_{0},t_{1},t_{2}$, $\{\Delta V_{tot,k}\}$
@@
4:   $\{\{\bm{\mu}_{k}(t)\},\{\bm{\mu}_{m,k}(t)\},\{\Sigma_{k}(t)\},\{\Sigma_{m,k}(t)\}\}\leftarrow$ Propagate 13 with $\{\Delta V_{safe,k}\}$.
5:   if 14b and 14c are satisfied then return $\{\Delta V_{safe,k}\}$.
@@
7:   Solve 14 with linearized version of 14b and 14c $\forall t\in\{t_{c}\}$.
@@
The optimization problem in Eq. 14 is a non-convex semi-infinite problem (SIP). To solve it efficiently, an iterative approach that is commonly used to solve SIPs is employed whereby affine inequality constraints are only applied at discrete time points at which the semi-infinite constraint is violated [40]. These time points are selected by finding the times at which the worst violation of the inequality constraints occurs. This recursive procedure is outlined in Alg. 1. In practice, the authors do not find it necessary to augment the non-convex constraints with slack or introduce a trust region for robust convergence. The CBF-SIP should be recursively feasible so long as the covariance at the next maneuver time does not become larger than the propagated covariance from the previous maneuver time and the estimated relative state error does not change drastically from one iteration to the next.
Typically, a maneuver profile is already free of collision risk, in which case the CBF returns the same input profile without running any optimization routine. This can be contrasted to the SCP-DRIFT and SCP-QPRIT techniques, which do not exploit the rarity of collision events. Moreover, leveraging the SIP reduction technique means that each optimization problem involves very few constraints, thereby improving scaling with respect to the number of agents. The CBF-SIP, however, does not necessarily provide guarantees of stability when combined with absolute and relative maneuver algorithms. In the case of very tight formations, where collision risk is higher, the CBF-SIP may result in greater dispersions and higher $\Delta V$ expenditure. For this reason, the CBF-SIP can be thought of as an additional form of (safe) disturbance to the formation that has magnitude proportional to the tightness of the formation.

### v2

@@ 3.3 Local Toroidal Coordinates About HFEM Trajectories
Local toroidal coordinates (LTCs) offer a convenient way to visualize, represent, and encode equilibrium in relative orbital geometries in restricted three-body problem orbits [elliott2022describing, elliott2022phd]. As a result of the procedure used to identify finite-duration QPROs about HFEM trajectories, it is possible to leverage LTCs for the duration of the QPRO despite the fact that the underlying reference is not periodic. Let $e_{r}+e_{i}i$ represent the quasi-unimodular eigenvalues and $\bm{w}_{r,0}+\bm{w}_{i,0}i$ represent their associated eigenvectors determined from a HFEM STM $\Phi(t_{1},t_{0})$. The mapping of these eigenvectors to any point in time along the reference trajectory is given by $\bm{w}_{r}(t)=\Phi(t,t_{0})\bm{w}_{r,0}$ and $\bm{w}_{i}(t)=\Phi(t,t_{0})\bm{w}_{i,0}$. Let the eigenvectors be decomposed into position and velocity components with $\bm{w}_{r}=\begin{bmatrix}\bm{r}_{r}^{\intercal},\bm{v}_{r}^{\intercal}\end{bmatrix}^{\intercal}$ and $\bm{w}_{i}=\begin{bmatrix}\bm{r}_{i}^{\intercal},\bm{v}_{i}^{\intercal}\end{bmatrix}^{\intercal}$. Local toroidal coordinates are defined as $\bm{\zeta}\triangleq\begin{bmatrix}\bm{\zeta}_{r}^{\intercal},\bm{\zeta}_{v}^{\intercal}\end{bmatrix}^{\intercal}$, $\bm{\zeta}_{r}\triangleq[\alpha,\beta,h]^{\intercal}$, $\bm{\zeta}_{v}\triangleq[\dot{\alpha},\dot{\beta},\dot{h}]^{\intercal}$ and can be related to a relative position and velocity with respect to the reference with
[…]
If the LTCs of a spacecraft are such that $h=0$ and $\bm{\zeta}_{v}=\bm{0}$, then the point lies on the surface of a locally invariant torus. Moreover, $\alpha$ and $\beta$ describe the location on the invariant circle, as is visualized in Fig. 6.
Figure 6: Formation configuration on an osculating invariant circle. Spacecraft can be separated by phase angle ($\theta_{1}$) or radius ($K$), and equivalently represented in local toroidal coordinates ($\alpha$$\beta$).
@@ 4.1 Absolute Control
To compute absolute station-keeping maneuvers that keep the formation in the vicinity of a reference, this work makes use of an MPC scheme similar to the one proposed in Ref. shimane2025output. However, the scheme proposed here instead computes just one maneuver (instead of two, as is done in Ref. shimane2025output) that targets a position ellipsoid at perilune crossing six revolutions downstream along the reference. This is akin to the popularly employed x-axis crossing scheme [davis2022orbit, Guzzetti2017], which involves computing one maneuver to target the state of the reference trajectory at six revolutions downstream, which has been found to produce favorable $\Delta V$ expenditures. Given a reference trajectory $\bm{\phi}(t)$, an initial condition $\bm{x}_{0}$ and initial time $t_{0}$, the $\Delta V_{abs}$ is solved for with
@@
$\|[I\bm{0}](\bm{x}_{f}-\bm{\phi}(t_{f}))\|_{2}\leq\epsilon,$
where $B=\begin{bmatrix}\bm{0}&I\end{bmatrix}^{\intercal}$, $t_{f}$ is the time corresponding to perilune crossing at six revolutions downstream from the current time, and $\bm{f}$ is a function representing the HFEM equations of motion of the system. Given the nonlinear nature of the optimization problem, it is solved using sequential convex programming (SCP) by enforcing a trust region about the solution variables, augmenting the dynamics constraint with a slack variable, and penalizing the $l_{2}$ norm of the slack in the objective. The only solution variables in the SCP problem are the initial condition dynamic slack variable $\bm{\nu}\in\mathbb{R}^{6}$, which is augmented to the objective with an $l_{2}$-norm penalty, and the $\Delta V_{abs}\in\mathbb{R}^{3}$, for which a trust region of 1 cm/s is enforced. The final position excursion ellipsoid is set to $\epsilon=$ 25 km. The optimization routine should be theoretically always feasible [shimane2025output]; and an appropriately applied SCP solution routine robustly and reliably converges [mao2019successiveconvexificationsuperlinearlyconvergent].
@@
Fig. 8 shows the performance of the employed absolute station-keeping scheme about the reference trajectory in Fig. 1 simulated with a ground truth including solar radiation pressure and with realistic navigation and control errors. Commanded maneuvers with magnitude below 0.15 cm/s are not executed, in line with existing studies of absolute station-keeping [Guzzetti2017, davis2022orbit].
@@ 4.2 Relative Control
Given the inadequacies of lower fidelity models in relative control tasks over long durations, relative dynamical motion is generated directly about the reference HFEM trajectory to more accurately represent the relative motion. Since state transition matrices are used to compute the HFEM trajectory and the HFEM QPROs, this relative motion is readily available and could even be stored upon or uplinked to a spacecraft to allow for autonomous relative control. Since dispersions from a reference relative trajectory are typically small in a station-keeping scenario, a simple and effective procedure for computing corrective impulsive maneuvers is to simply target a reference position along the QPRO at some later time [elliott2022phd]. Concretely, to compute a relative maneuver $\Delta V_{rel}$ to deliver the $k$th spacecraft to a desired position $\delta\bm{r}_{f,k}$ (e.g., a reference position along a QPRO) at time $t_{f}$ from some initial state $\delta\bm{x}_{0,k}$ at $t_{0}$ simply requires solving
@@
Here, $\Phi_{rv}$ and $\Phi_{rx}$ denote the upper right block and right block of the STM respectively. Under ideal scenarios, the relative state will converge to the reference equilibrium in just two impulses. It is efficient to compute, scalable to multiple agents, and independent of the choice of dynamics. Though the control law makes no consideration for fuel expenditure, relative motion is typically so sensitive to changes in velocity in the station-keeping scenario that the commanded $\Delta V$ is usually only on the order of millimeters per second.
@@ 4.3 Safety Concept
A particular challenge in cislunar formation control is that there exists no known analytical condition which guarantees passive safety of spacecraft in a formation. In Keplerian orbits, passive safety is guaranteed by maintaining perpendicularity between spacecraft relative positions and the direction of the unstable mode. An analytical condition, known as eie-i vector separation, has been derived, which when satisfied provides long duration guarantees of safe spacecraft separation under perturbations [damico2006proximity]. In contrast, there does not exist an equally simple and strong notion of passive safety for cislunar orbits. In fact, from a QPRO equilibrium with kilometer level separation, a change in velocity on the order of millimeters per second can deliver a spacecraft to an unsafe configuration in just the span of weeks. In the case in which a spacecraft misses a maneuver, collisions can potentially occur within days or hours. This phenomenon is attributable to the existence of stable and unstable modes of the time-varying eigensystem, whose directions are constantly changing over the span of an orbit.
@@
Collision risk is usually not considered in studies of cislunar formation flight, though solutions that have been proposed include passively-safe sequential convex programming (SCP-DRIFT), whereby nonlinear optimization routines determine a sequence of maneuvers that guarantees passive safety, and maneuver planning leveraging local-toroidal elements (SCP-QPRIT), which allows for elegant suppression of stable and unstable modes [takubo2025safeoptimalnspacecraftswarm]. SCP is computationally expensive and scales poorly with agent count, which makes it impractical for long-duration station-keeping. Additionally, LTCs, which provide weaker guarantees of safety, are not amenable to the inclusion of absolute station-keeping inputs, which are typically spanned by unstable or stable modes.
@@
The safety concept proposed in this work leverages an optimization-based control barrier function (CBF) to provide safety guarantees to the formation. This approach takes advantage of the rarity of collision events, provides recursive safety, and allows for decoupled computation of the absolute and relative control inputs. CBFs typically involve using optimization protocols to “filter” a nominal control input to be safe [ames2019controlbarrierfunctionstheory]. Though not often applied to spacecraft and typically considered for systems with continuous control, CBFs and similar procedures have been employed for impulsively and continuously controlled spacecraft in Refs. Stephenson2025a and vanwijkrlcbf for ensuring safety in reinforcement learning routines. The variant of CBF proposed here involves enforcing stochastic keep out zone constraints in a similar manner to passively safe optimal control schemes [guffanti2023passively] and represents one of the first applications of safety filtering to cislunar formation flight.
@@
Let $t_{CO,max}$ and $t_{CO,min}$ be the maximum and minimum cutoff time to a spacecraft’s propulsion system following a random missed maneuver event at time $t_{0}$. It is assumed that following some cutoff time in this interval, the propulsion system will be guaranteed to function at the next available control time. Under these conditions, it is possible to “filter” the desired control input of all agents at $t_{0}$ by enforcing keep out zones over the evolution of the agents over the time interval. Particularly, the evolution of the agents subject to commanded control inputs $\{\Delta V_{tot,k}=\Delta V_{abs}+\Delta V_{rel,k}\}_{k=1}^{M}$ subject to some initial system state provided by the navigation filter, $\bm{X}_{0}\sim\mathcal{N}(\bm{\mu}_{0},\Sigma_{0})$, where $\bm{X}_{0}=[\bm{x}_{1}^{\intercal},...,\bm{x}_{M}^{\intercal}]^{\intercal}$, $\mathbb{E}(\bm{x}_{k})=\bm{\mu_{k}}$, $\mathbb{E}((\bm{x}_{k}-\bm{\mu}_{k})(\bm{x}_{j}-\bm{\mu}_{j})^{\intercal})=\Sigma_{kj}$ can be propagated as a distribution with
$\begin{gathered}\bm{\mu}_{k}(t)=\bm{\mu}_{k,0}+B\Delta V_{k}+\int_{t_{0}}^{t}\bm{f}(\bm{x}(\tau),\tau)d\tau,\\ \Sigma_{kj}(t)=\Phi_{k}(t,t_{0})(\Sigma_{kj}+\delta_{kj}BG_{exe,k}G^{\intercal}_{exe,j}B^{\intercal})\Phi^{\intercal}_{j}(t,t_{0})+\delta_{kj}\int_{t_{0}}^{t}\Phi_{k}(t,\tau)G_{rel}G_{rel}^{\intercal}\Phi_{k}^{\intercal}(t,\tau)d\tau\end{gathered}$ (13)
@@
Here, $G_{exe,k}$ is an estimate of the maneuver execution error associated with $\Delta V_{k}$ as determined from Gate’s model [Gates1963], $G_{rel}$ is the covariance associated with the perturbative acceleration to the relative states, $\Phi_{k}(t,t_{0})$ is the STM produced about the mean trajectory of the $k$th agent, $\bm{\mu}_{k}(t)$, and $\delta_{kj}$ is the Kronecker delta. The above distribution propagation can be simplified by approximating each state mean and covariance as linear with respect to the absolute station-keeping trajectory, whose state and STM is already available. The state can also be propagated in the case of missed maneuvers $\bm{x}_{k,m}\sim\mathcal{N}(\bm{\mu}_{k,m},\Sigma_{k,m})$ simply by setting $\Delta V_{k}=0$. To enforce passive safety, it is necessary to ensure that the distributions of each agent are non-overlapping to within some probability for all time until the next available control input. The approach taken here is to enforce that any two spacecraft remain a distance $D_{min}$ apart under $3\sigma$ certainty [guffanti2023passively], which can be attained by computing a small modification to the nominal control input through solving a nonconvex optimization problem. Let $t_{1}$ denote the first control input time after $t_{0}+t_{CO,max}$ and $t_{2}$ denote the first control input time after $t_{1}+t_{CO,max}$. Furthermore, let $\bm{r}_{kj}(t)=R(\bm{x}_{k}(t)-\bm{x}_{j}(t))$ be the relative position between spacecraft $k$ and $j$, produced by setting $R=[I_{3\times 3}\ \bm{0}_{3\times 3}]$, which is normally distributed according to $\bm{r}_{kj}(t)\sim\mathcal{N}(\bm{p}_{kj}(t),P_{kj}(t))$ where $\bm{p}_{kj}(t)=R(\bm{\mu}_{k}(t)-\bm{\mu}_{j}(t))$ and $P_{kj}(t)=R(\Sigma_{k}(t)-\Sigma_{kj}(t)-\Sigma_{jk}(t)+\Sigma_{j}(t))R^{\intercal}$. Note that the covariance of the relative position depends not only the covariances of the states $k$th and $j$th spacecraft, but also the covariance of the states of the states $k$th and $j$th spacecraft with respect to each other. Then, $\{\Delta V_{safe,k}\}_{k=1}^{M}$ can be solved for from a desired $\{\Delta V_{tot,k}\}_{k=1}^{M}$ by iteratively solving the problem
$\min_{\{\Delta V_{safe,k}\}_{k=1}^{M}}\quad$ $\sum_{k=1}^{M}\|\Delta V_{safe,k}-\Delta V_{tot,k}\|$ (14a)
s.t. $\bm{p}^{\intercal}_{kj}(t)\bm{p}^{(q-1)}_{kj}(t)-3\|P^{\frac{\intercal}{2}}_{kj}(t)\bm{p}^{(q-1)}_{kj}(t)\|\geq D_{min}\|\bm{p}^{(q-1)}_{kj}(t)\|\ \forall t\in[t_{0},t_{2}],\ \forall k<j$ (14b)
$\bm{p}^{\intercal}_{kj,m}(t)\bm{p}^{(q-1)}_{kj,m}(t)-3\|P^{\frac{\intercal}{2}}_{kj}(t)\bm{p}^{(q-1)}_{kj,m}(t)\|\geq D_{min}\|\bm{p}^{(q-1)}_{kj,m}(t)\|\ \forall t\in[t_{0},t_{1}],\ \forall k\neq j$ (14c)
Here, $\bm{p}^{(q-1)}_{kj}(t)$ and $\bm{p}^{(q-1)}_{kj,m}(t)$ denote the relative positions at time $t$ as computed from the (q1)(q-1)th iteration of the scheme. When the scheme is converged, i.e., $\bm{p}^{(q-1)}_{kj}(t)=\bm{p}^{(q)}_{kj}(t)$ and $\bm{p}^{(q-1)}_{kj,m}(t)=\bm{p}^{(q)}_{kj,m}(t)$, the stochastic keep out zones are enforced over the entire duration [guffanti2023passively]. Note that constraint Eq. 14b enforces stochastic keep out zones until the second available maneuver time ($t_{2}$) between all successfully maneuvered trajectories whereas constraint Eq. 14c enforces stochastic keep out zones until the first available maneuver time ($t_{1}$) between all possibly failed maneuvered trajectories. The former constraint ensures that in the case of two spacecraft missing maneuvers at the next maneuver time ($t_{1}$), the spacecraft will remain collision free until the following maneuver time ($t_{2}$), at which time both spacecraft will be capable of maneuvering under the previously made assumption of a maximum propulsion cutoff time. These features of the CBF endow the system with recursive safety under the assumption that a spacecraft will be able to execute a maneuver following the cutoff period of an MME.
@@
Input $\{\bm{\mu}_{0}\}$, $\{\Sigma_{0}\}$, $t_{0},t_{1},t_{2}$, $\{\Delta V_{tot,k}\}$
@@
4:   $\{\{\bm{\mu}_{k}(t)\},\{\bm{\mu}_{k,m}(t)\},\{\Sigma_{kj}(t)\}\}\leftarrow$ Propagate 13 with $\{\Delta V_{safe,k}\}$.
5:   if 14b and 14c are satisfied abouth current reference then return $\{\Delta V_{safe,k}\}$.
@@
7:   $\{\Delta V_{safe,k}\}\leftarrow$ Solve 14 with linearized version of 14b and 14c $\forall t\in\{t_{c}\}$.
@@
The optimization problem in Eq. 14 is a non-convex semi-infinite problem (SIP). To solve it efficiently, an iterative approach that is commonly used to solve SIPs is employed whereby affine inequality constraints are only applied at discrete time points at which the semi-infinite constraint is violated [Shapiro01022009]. These time points are selected by finding the times at which the worst violation of the inequality constraints occurs. This recursive procedure is outlined in Alg. 1. In practice, the authors do not find it necessary to augment the non-convex constraints with slack or introduce a trust region for robust convergence. The CBF-SIP should be recursively feasible so long as the covariance at the next maneuver time does not become larger than the predicted covariance from the previous maneuver time and the estimated relative state error does not change drastically from one iteration to the next.
Typically, a maneuver profile is already free of collision risk, in which case the CBF returns the same input profile without running any optimization routine. This can be contrasted to the SCP-DRIFT and SCP-QPRIT techniques, which do not exploit the rarity of collision events for efficient solution procedures. Moreover, leveraging the SIP reduction technique means that each optimization problem involves very few constraints, thereby improving scalability with respect to the number of agents. The CBF-SIP, however, does not necessarily provide guarantees of stability when combined with absolute and relative maneuver algorithms. In the case of very tight formations, where collision risk is higher, the CBF-SIP may result in greater dispersions and higher $\Delta V$ expenditure. For this reason, the CBF-SIP can be thought of as an additional form of (safe) disturbance to the formation that has magnitude proportional to the tightness of the formation.